25 &27

In Problems 15-28 find the general solution of the given higher-order differential equation. 15 y – 4y – 5y = 0 16. y – y

d²x d²x an de dt2 4x = 0 21. y + 3y + 3y + y = 0 22. y – 6y + 12y – 8y = 0 (23. + y + y = 0 24. y(4) - 2y + y = 0 dº
Answer #1

7: 25 To dy peady tayo dut dae Its auxillomyequution is 1674 +24 m2 +9=0 (4 m2)2+ 8?+ 2x4mn 3 =0.. 14 m2 +3.92=0 4m²+3= 6 m²I hope you like this.